# Discovering universal temperature regulation dynamics in animals

### v2

@@
Discovering universal temperature regulation dynamics in animals
Cody E. FitzGerald Department of Engineering Sciences and Applied Mathematics, Northwestern University, Evanston, IL 60208, USA cody.fitzgerald@northwestern.edu Andrew J. Engedal Lewis-Sigler Institute of Integrative Genomics, Princeton University, Princeton, NJ 08540, USA Niall M. Mangan Department of Engineering Sciences and Applied Mathematics, Northwestern University, Evanston, IL 60208, USA NSF-Simons National Institute for Theory and Mathematics in Biology, Chicago, IL 60611, USA
(August 25, 2025)
1 Abstract
Hibernation is an adaptation to extreme environmental seasonality that has been studied for almost 200 years, but our understanding of the underlying physiological system remains lacking due to the partially observed nature of the system. During hibernation, small mammals, such as the Arctic ground squirrel, exhibit dramatic oscillations in body temperature, typically one of the only physiological states measured, of up to 40 C. These spikes are known as interbout arousals and typically occur 10-20 times throughout hibernation. The physiological process that drives interbout arousals is unknown, but two distinct macro-scale mechanisms have been hypothesized. Using model selection for partially observed systems and classical dynamical systems theory, we are able to differentiate between these two hypotheses using only body temperature data recorded from a free-ranging Arctic ground squirrel, and show that our model can capture the broad features of the observed seasonal physiological transitions. We then modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, a shrew, and a bear, which also dynamically modulate body temperature. Our results suggest that a universal, environmentally sensitive core regulator could control body temperature across a diverse range of species—a restructuring of our current understanding of the physiological organization across species. While the findings presented here are applicable to thermophysiology, the general modeling procedure is applicable to time series data collected from partially observed biological, chemical, physical, mechanical, and cosmic systems for which the goal is to elucidate the underlying mechanism or control structure.
@@ 2 Introduction
Scientists have studied hibernation for nearly two hundred years [1] in the hopes of gaining insight into this complex physiological system and to extend our own physiology. Hibernation research has a wide range of practical applications to human health, including cardiac arrest, organ preservation, and longevity. Inducing hibernation-like states into the range of human physiology is also under study, as this may help realize our species’ goals of long-distance human space travel [2, 3]. Interest in hibernation and dormancy states has transcended science and also fills our literature. Proust writes of “…the comfort of reclusion, the poetry of hibernation…” in A Swan’s Way. A more recent novel, My Year of Rest and Relaxation, follows a New York City woman who entered a months-long hibernation-like state. But for all this sustained scientific and broader interest, our understanding of dynamic temperature regulation (known as the thermophysiological system) that facilitates hibernation remains limited, in part because few physiological states of the system can be observed experimentally. Typically, body temperature is the only physiological measurement that can be recorded from free-ranging organisms, though surely other states of the physiological system must contribute to thermoregulation. In applied mathematics, this is known as a partially observed system, and inferring models for this class of system is technically challenging and an active area of research [4, 5]. Here, we combine body temperature measurements with automated model discovery and classical dynamical systems theory to develop a mechanistic understanding of small mammalian thermophysiology and uncover a universal mechanism of thermophysiology across species.
@@
We begin our search for a universal mechanism by studying the Arctic ground squirrel, a species that has been well-studied and exhibits seasonal physiological changes representative of many small mammalian hibernating species. The Arctic ground squirrel lives in the extremes of the high North American Arctic, and it experiences dramatic seasonal physiological changes. In a typical year, an Arctic ground squirrel remains above ground for approximately four months and hibernates underground for about eight months. During the summer months, the Arctic ground squirrel is homeothermic, meaning its body temperature hovers around 37 C with minimal daily fluctuations caused by activity. A seasonal thermoregulatory transition occurs as summer merges into fall, and the Arctic ground squirrel enters hibernation (Figure 1). At the start of hibernation, the Arctic ground squirrel makes a thermoregulatory transition from homeothermy to poikilothermy, a thermoregulatory regime in which body temperature is held near or below 0 C, with occasional spikes back to 37 C known as “interbout arousals.” During this period, the body temperature of the Arctic ground squirrel can be as low as -2.9 C. The organism can lower its body temperature below the freezing point through a process called super-cooling [6]. Toward the end of hibernation, the Arctic ground squirrel makes the reciprocal thermoregulatory switch from poikilothermy to homeothermy, and the annual cycle continues. An open challenge is determining what underlying mechanisms control these dramatic changes in body temperature.
@@
Relatively few mathematical models of hibernation and thermophysiology exist, and those that do exist tend to be either relatively complex and thus difficult to validate against the existing data or statistical in nature. Hampton and Andrews put forward a mechanistic model of small mammalian hibernation [10]. Their model puts forward a mechanism that triggers interbout arousals and can be viewed as a version of the Hourglass and Threshold Hypothesis. However, the model formulation is very high-dimensional and cannot be validated using the available experimental data. The authors also developed a complex physiological model to infer metabolic rate from other physiological data [11]. Others have used statistical frameworks to examine features of hibernation. Oro and Freixas used statistical modeling to show that air temperature drives “flickering,” a behavior that occurs as the organism transitions between homeothermy (constant body temperature) and poikilothermy (spiking body temperature) [12]. A different study used Generalized Harmonic Analysis to investigate interbout arousals in small hibernators [13]. We seek to find the right level of mathematical description for hibernation physiology that balances interpretability and preserves the ability to validate the model against the existing data. This is challenging due to natural and experimental limitations intrinsic to the system.
@@
Hibernation is a multi-scale physiological adaptation that occurs on the year scale. For this reason, a molecular description of the biological processes that underpin hibernation will be exceedingly difficult to find. It is unlikely we will ever have complete and comprehensive data sets of all relevant neurophysiological subsystems involved in hibernation due to the costs of studying hibernators from both a financial and animal welfare perspective. Most of the data collected from hibernation experiments occurs at the physiological scale rather than the molecular scale. We grappled with the right size and type of model to use in describing this system, and ultimately found a model that is roughly as mechanistic the Fitzhugh-Nagamo model [14, 15], which describes neuronal spiking dynamics; the NPZ model [16], which describes the dynamics of nutrients, phytoplankton and zooplankton; or the Brusselator model [17, 18, 19] which describes the Belousov-Zhabotinsky chemical reaction [20]. The goal of these models, and perhaps all models, is to describe the essence of the underlying process in an interpretable manner without being completely overrun by the complexity of the system or the curse of dimensionality. Balancing model simplicity, which allows for interpretability, with complexity, which allows for greater realism, is a fine line. The NPZ model is a good example of a model that exhibits this balance. It captures a complex, multi-scale ecological system using just three coupled equations, and can be used to probe how changes or perturbations to one state influence the others, which would be challenging to understand both without any model and with too complex of a model. Here, we attempt to find a similar modeling balance. We discover a macro-scale model with interpretable parameters that have physiological meaning, that is complex enough to capture the main features of the data, that is simple enough to be fit to the existing physiological data, and can be used to guide the search for microscale mechanisms that underpin hibernation. Using our discovered model, we provide new insights into hibernation physiology, seasonal physiological transitions, and the possible conserved physiological organization across species.
@@ 3.1 Model selection, robustness testing, and validation
To find a thermophysiological model for the Arctic ground squirrel, we employ a recent sparse model selection algorithm designed for partially observed systems [4]. In our system, only one variable, body temperature, is measured (Figure 2A). To determine the intrinsic dimensionality, we statistically analyze time-delay embeddings of body temperature data and find that the system is two-dimensional (Figure 2A and Section 6.2) [28]. Thus, we seek to find a 2-D system of Ordinary Differential Equations (ODE) comprised of monomial terms where the states are the measured temperature, $x$, and a hidden state, $y$, which represents the dynamics of an unknown physiological process. This allows us to find an effective macro-scale physiological model describing the dynamics of coupling between temperature and some unknown internal control process. Sparse model selection evaluates a library of possible terms, ${\bf\Theta}(x,y)$, and composes ODE models by identifying a parsimonious set of terms that match observations of a subset of the system variables (Figure 2A-B). The search is performed by iteratively estimating the coefficients in the library using variational annealing and applying a threshold as a sparsity penalty. Iterative thresholding on the magnitude of the coefficients for each model term selects terms with large coefficients and sets small coefficients to zero in the next iteration. Of the models that emerge from varying sparsity and increasing model enforcement during optimization, those along the Pareto front balance simplicity and fit to the data, so we select the best fitting eight, nine, and ten active terms for further investigation (Figure 2B). [^1] For specific parameters the seven-term model exhibited a limit cycle, but cycling was not robust to small changes to any of the optimized parameters. These models form a nested family, and the algorithm selects variations on the same model. Our three models of interest can be written in the form
@@
To enable the models of interest to exhibit limit cycles with a finite period, we introduced a constant leak term, $\theta_{0}$, to the body temperature equation (see Equations (3)-(4)). We found that several low-order terms could be introduced to regulate the period of the limit cycle, but we chose to modify the data-driven models in the simplest way possible.  [^3] Notably, there exists a transformation from our models to Líenard equations when a linear term in $y$ is added to the body temperature equation [31]. With this simple addition, the models of interest generate stable limit cycles with finite periods. To show that the limit cycle approaches an infinite period with zero leak, we introduced the leak term and simulated the Equations (3)-(4) (with $\theta_{C}=\theta_{D}=0$) as $\theta_{0}\rightarrow 0$. As expected, this resulted in limit cycle behavior with increasingly long periods, as the phase plane velocity near the origin goes to zero as $\theta_{0}\rightarrow 0$. Next, we perform additional robustness testing to determine which structural terms are necessary among the nested family of models.
@@ 3.2 Term robustness testing
Here, we test whether the remaining terms are necessary to produce the observed behavior based on overall term size. This is equivalent to keeping only terms that contribute significantly to the dynamics, consistent with dominant balance [32, 33, 34]. By forward simulating Equations (3)-(4) with $\theta_{C}=\theta_{D}=0$ and examining the term sizes of each component, we find that two additional terms, $\theta_{A}y^{3}$ and $\theta_{B}xy^{2}$, are negligible across the time-series of a spike (see Figure 2C). The terms $\theta_{0}$, $\theta_{5}x^{2}$, and $\theta_{6}y^{2}$ are excluded from this analysis as they are required to generate finite period limit cycle behavior. We now have a larger nested family of models of the form
@@
where the terms with coefficients $\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5},\theta_{6}$ (in black), define the core model and the additional terms with coefficients, $\theta_{A},\theta_{B},\theta_{C},\theta_{D}$ (blue and pink, Figure 2 D), are non-essential higher-complexity terms. Next, we validate our models using the $l^{2}$ norm, information criteria, and dynamic time warping.
@@ 3.3 Model validation
To quantitatively evaluate the fit of the model ensemble data to the data, we re-estimate the parameters of each model and compare against the training data (Figure 3A). We refer to the models by the number of active terms (not including the introduced leak term, $\theta_{0}$) and the subscript of the parameters of higher complexity terms that the model contains (see Eqs. (3)-(4)). For example, Model 8AB refers to Eqs. (3)-(4) with $\theta_{A}\neq 0$, $\theta_{B}\neq 0$, and $\theta_{C}=\theta_{D}=0$. A complication in parameter estimation and validation for these models is that the interbout arousals of the Arctic ground squirrel are slightly aperiodic, perhaps due to small fluctuations in the decay rate of the unknown metabolite [8]. To generate a good fit, the forward simulation must match both the height of each body temperature spike and its timing, but a periodic spike will always be misaligned with the aperiodic data. To compensate, the peaks flatten to match most of the non-spiking data, which is near zero. To avoid forcing the data to match multiple periods, we fit all 16 models to the first two spikes of the training data (Figure 3A). We then evaluate the fit of the models using the $l^{2}$ norm and information criteria on the first two spikes of the training data, and validate the models on out-of-sample data using dynamic time warping.
@@ 3.4 Discovered model is consistent with Hourglass and Threshold Hypothesis
The dynamics of the hidden physiological state, $y$, of Model 6 appear with a constant period rather than a period that varies with body temperature (Figure 3C). This supports the Hourglass and Threshold Hypothesis over the Torpor Arousal Clock Hypothesis. The ability to differentiate between the two hypotheses is not dependent on the parameterization and structure of the model; the dimensionality of the thermophysiological system alone supports a constant period and the depleted metabolite hypothesis. In two dimensions, planar ODE systems cannot give rise to non-trivial solutions in which $x(t)$ and $y(t)$ have different periods. However, we can study the identified model’s structure, parameterization, and symmetries to extract additional physiological implications.
Directly relating the hidden state $y(t)$ to a particular metabolite or physiological state is impossible without direct measurements. However, we can computationally explore possible mechanisms and hypothesize about the minimal mechanism of regulation. In the simplest case, we can imagine the hidden state as a molecule that is depleting over time as envisioned by Ruf, Giroud, and Geiser [8] (see Figure 3C top). However, by introducing a symmetry transformation into the hidden state of the form $\tilde{y}=-y+\psi$, we can generate dynamics that become negative (see Figure 3C bottom) or correspond to an accumulating, potentially toxic molecule [8]. If we take $\psi=0.1$, then the scaling symmetry may correspond to a physiological scenario where the threshold acts on the difference between multiple molecules rather than an individual entity (See Figure 3 C bottom). If we take $\psi=0.75$, then $\tilde{y}(t)$ stays completely positive and inverts the dynamics of $y(t)$ (top, Figure 3C). This parameterization of the scaling symmetry reflects a fact mentioned in [8]: there is no way to distinguish between an accumulating metabolite and a depleting metabolite as the driver of interbout arousals. This class of symmetry transformations does not affect the fit to the measured state, so we cannot distinguish which of these model variants and corresponding metabolic scenarios are correct. Whether representing a single molecule or interactions between multiple molecules, the effective state acts as an Hourglass and Threshold-type mechanism.

### v4

# Discovering a low-dimensional temperature control architecture across animals

@@
Abstract
Hibernation is an adaptation to extreme environmental seasonality that has been studied for almost 200 years, but our understanding of the underlying physiological system remains lacking due to the partially observed nature of the system. During hibernation, small mammals, such as the Arctic ground squirrel, exhibit dramatic oscillations in body temperature, typically one of the only physiological states measured, of up to 40 C. These spikes are known as interbout arousals and typically occur 10-20 times throughout hibernation. The physiological process that drives interbout arousals is unknown, but two distinct macro-scale mechanisms have been hypothesized. Using model selection for partially observed systems and classical dynamical systems theory, we are able to differentiate between these two hypotheses using only body temperature data recorded from a free-ranging Arctic ground squirrel, and show that our model can capture the broad features of the observed seasonal physiological transitions. We then modify our discovered physiological model of Arctic ground squirrel to include internally-encoded environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, a shrew, and a bear, which also dynamically modulate body temperature. Our results suggest that a low-dimensional, environmentally sensitive core regulator could control body temperature across a diverse range of species—a new understanding of the physiological organization across species. While the findings presented here are applicable to thermophysiology, the general modeling procedure is applicable to time series data collected from partially observed biological, chemical, physical, mechanical, and cosmic systems for which the goal is to elucidate the underlying mechanism or control structure.
keywords:
hibernation, physiology, dynamical systems, model selection
@@ 1 Introduction
Scientists have studied hibernation for nearly two hundred years [1] in the hopes of gaining insight into this complex physiological system and to extend our own physiology. Hibernation research has a wide range of practical applications to human health, including cardiac arrest, organ preservation, and longevity. Inducing hibernation-like states into the range of human physiology is also under study, as this may help realize our species’ goals of long-distance human space travel [2, 3]. Interest in hibernation and dormancy states has transcended science and also fills our literature. Proust writes of “…the comfort of reclusion, the poetry of hibernation…” in A Swan’s Way. A more recent novel, My Year of Rest and Relaxation, follows a New York City woman who entered a months-long hibernation-like state. But for all this sustained scientific and broader interest, our understanding of dynamic temperature regulation (known as the thermophysiological system) that facilitates hibernation remains limited, in part because few physiological states of the system can be observed experimentally. Typically, body temperature is the only physiological measurement that can be recorded from free-ranging organisms, though surely other states of the physiological system must contribute to thermoregulation. In applied mathematics, this is known as a partially observed system, and inferring models for this class of system is technically challenging and an active area of research [4, 5]. Here, we combine body temperature measurements with automated model discovery and classical dynamical systems theory to develop a mechanistic understanding of small mammalian thermophysiology. We then show that our discovered model acts as a low-dimensional control architecture capable of describing observed temperature dynamics across multiple species.
@@
We begin our search for a low-dimensional temperature controller by studying the Arctic ground squirrel, a species that has been well-studied and exhibits seasonal physiological changes representative of many small mammalian hibernating species. The Arctic ground squirrel lives in the extremes of the high North American Arctic, and it experiences dramatic seasonal physiological changes. In a typical year, an Arctic ground squirrel remains above ground for approximately four months and hibernates underground for about eight months. During the summer months, the Arctic ground squirrel is homeothermic, meaning its body temperature hovers around 37 C with minimal daily fluctuations caused by activity. A seasonal thermoregulatory transition occurs as summer merges into fall, and the Arctic ground squirrel enters hibernation (Figure 1). At the start of hibernation, the Arctic ground squirrel makes a thermoregulatory transition from homeothermy to the hibernation thermoregulatory regime in which body temperature alternates between homeothermy near or below 0 C and occasional returns to euthermic conditions near 37 C. These returns to euthermic conditions are known in the biological literature as “interbout arousals.” During hibernation, the body temperature of the Arctic ground squirrel can be as low as -2.9 C. The organism can lower its body temperature below the freezing point through a process called super-cooling [6]. Toward the end of hibernation, the Arctic ground squirrel makes the reciprocal thermoregulatory switch from the hibernation thermoregulatory regime to homeothermy, and the annual cycle continues. An open challenge is determining what underlying mechanisms control these dramatic changes in body temperature.
@@
Relatively few mathematical models of hibernation and thermophysiology exist, and those that do exist tend to be either relatively complex and thus difficult to validate against the existing data or statistical in nature. Hampton and Andrews put forward a mechanistic model of small mammalian hibernation [10]. Their model puts forward a mechanism that triggers interbout arousals and can be viewed as a version of the Hourglass and Threshold Hypothesis. However, the model formulation is very high-dimensional and cannot be validated using the available experimental data. The authors also developed a complex physiological model to infer metabolic rate from other physiological data [11]. Others have used statistical frameworks to examine features of hibernation. Oro and Freixas used statistical modeling to show that air temperature drives “flickering,” a behavior that occurs as the organism transitions between homeothermy (constant body temperature) and the hibernation thermoregulatory regime (spiking body temperature) [12]. A different study used Generalized Harmonic Analysis to investigate interbout arousals in small hibernators [13]. We seek to find the right level of mathematical description for hibernation physiology that balances interpretability and preserves the ability to validate the model against the existing data. This is challenging due to natural and experimental limitations intrinsic to the system.
@@
Hibernation is a multi-scale physiological adaptation that occurs on the year scale. For this reason, a molecular description of the biological processes that underpin hibernation will be exceedingly difficult to find. It is unlikely we will ever have complete and comprehensive data sets of all relevant neurophysiological subsystems involved in hibernation due to the costs of studying hibernators from both a financial and animal welfare perspective. Most of the data collected from hibernation experiments occurs at the physiological scale rather than the molecular scale. We grappled with the right size and type of model to use in describing this system, and ultimately found a model that is roughly as mechanistic the Fitzhugh-Nagamo model [14, 15], which describes neuronal spiking dynamics; the NPZ model [16], which describes the dynamics of nutrients, phytoplankton and zooplankton; or the Brusselator model [17, 18, 19] which describes the Belousov-Zhabotinsky chemical reaction [20]. The goal of these models, and perhaps all models, is to describe the essence of the underlying process in an interpretable manner without being completely overrun by the complexity of the system or the curse of dimensionality. Balancing model simplicity, which allows for interpretability, with complexity, which allows for greater realism, is a fine line. The NPZ model is a good example of a model that exhibits this balance. It captures a complex, multi-scale ecological system using just three coupled equations, and can be used to probe how changes or perturbations to one state influence the others, which would be challenging to understand both without any model and with too complex a model. Here, we attempt to find a similar modeling balance. We discover a macro-scale model with interpretable parameters that have physiological meaning, that is complex enough to capture the main features of the data, that is simple enough to be fit to the existing physiological data, and can be used to guide the search for microscale mechanisms that underpin hibernation. Using our discovered model, we provide new insights into hibernation physiology, seasonal physiological transitions, and the possible conserved physiological organization across species.
@@ 2.1 Model selection, robustness testing, and validation
To find a thermophysiological model for the Arctic ground squirrel, we employ a recent sparse model selection algorithm designed for partially observed systems [4]. In our system, only one variable, body temperature, is measured (Figure 2A). To determine the intrinsic dimensionality, we statistically analyze time-delay embeddings of body temperature data and find that the system is two-dimensional (Figure 2A and Section 4.2) [28]. Thus, we seek to find a 2-D system of Ordinary Differential Equations (ODE) comprised of monomial terms where the states are the measured temperature, $x$, and a hidden state, $y$, which represents the dynamics of an unknown physiological process. This allows us to find an effective macro-scale physiological model describing the dynamics of coupling between temperature and some unknown internal control process. Sparse model selection evaluates a library of possible terms, $\boldsymbol{{\Theta}}(x,y)$, and composes ODE models by identifying a parsimonious set of terms that match observations of a subset of the system variables (Figure 2A-B). The search is performed by iteratively estimating the coefficients in the library using variational annealing and applying a threshold as a sparsity penalty. Iterative thresholding on the magnitude of the coefficients for each model term selects terms with large coefficients and sets small coefficients to zero in the next iteration. Of the models that emerge from varying sparsity and increasing model enforcement during optimization, those along the Pareto front balance simplicity and fit to the data, so we select the best fitting eight, nine, and ten active terms for further investigation (Figure 2B). [^1] For specific parameters the seven-term model exhibited a limit cycle, but cycling was not robust to small changes to any of the optimized parameters. These models form a nested family, and the algorithm selects variations on the same model. Our three models of interest can be written in the form
@@
To enable the models of interest to exhibit limit cycles with a finite period, we introduced a constant term representing a low, basal rate of heat production, $\theta_{0}$, to the body temperature equation (see Equation (3)). While metabolic rate and heat production are suppressed during hibernation, some small amount of heat is produced even during torpor to regulate the body temperature, $x$, and ensure the hibernator does not freeze [31, 32, 33]. We found that several low-order terms could be introduced to regulate the period of the limit cycle, but we chose to modify the data-driven models in the simplest way possible.  [^3] Notably, there exists a transformation from our models to Líenard equations when a linear term in $y$ is added to the body temperature equation [34]. With this simple addition, the models of interest generate stable limit cycles with finite periods. To show that the limit cycle approaches an infinite period with zero leak, we introduced the leak term and simulated the Equations (3)-(4) (with $\theta_{C}=\theta_{D}=0$) as $\theta_{0}\rightarrow 0$. As expected, this resulted in limit cycle behavior with increasingly long periods, as the phase plane velocity near the origin goes to zero as $\theta_{0}\rightarrow 0$. Next, we perform additional robustness testing to determine which structural terms are necessary among the nested family of models.
@@ 2.2 Term robustness testing
Here, we test whether the remaining terms are necessary to produce the observed behavior based on overall term size. This is equivalent to keeping only terms that contribute significantly to the dynamics, consistent with dominant balance [35, 36, 37]. By forward simulating Equations (3)-(4) with $\theta_{C}=\theta_{D}=0$ and examining the term sizes of each component, we find that two additional terms, $\theta_{A}y^{3}$ and $\theta_{B}xy^{2}$, are negligible across the time-series of a spike (see Figure 2C). The terms $\theta_{0}$, $\theta_{5}x^{2}$, and $\theta_{6}y^{2}$ are excluded from this analysis as they are required to generate finite period limit cycle behavior. In particular, $\theta_{5}x^{2}$ is needed for the existence of a limit cycle, otherwise there is no $x$-dependence in Eq. (4). The parameters $\theta_{0}$ and $\theta_{6}y^{2}$ help to ensure the limit cycle has a finite period by keeping the velocity in the phase plane non-zero near the origin. We now have a larger nested family of models of the form
@@
where the terms with coefficients $\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5},\theta_{6}$ (in black), define the core model and the additional terms with coefficients, $\theta_{A},\theta_{B},\theta_{C},\theta_{D}$ (blue and pink, Figure 2 D), are non-essential higher-complexity terms. Next, we validate our models using the $l^{2}$ norm, Akaike information criteria, and dynamic time warping.
@@ 2.3 Model validation
To quantitatively evaluate the fit of the model ensemble data to the data, we re-estimate the parameters of each model and compare against the training data (Figure 3A). We refer to the models by the number of active terms (not including the introduced leak term, $\theta_{0}$) and the subscript of the parameters of higher complexity terms that the model contains (see Eqs. (3)-(4)). For example, Model 8AB refers to Eqs. (3)-(4) with $\theta_{A}\neq 0$, $\theta_{B}\neq 0$, and $\theta_{C}=\theta_{D}=0$. A complication in parameter estimation and validation for these models is that the interbout arousals of the Arctic ground squirrel are slightly aperiodic, perhaps due to small fluctuations in the decay rate of the unknown metabolite [8]. To generate a good fit, the forward simulation must match both the height of each body temperature spike and its timing, but a periodic spike will always be misaligned with the aperiodic data. To compensate, the peaks flatten to match most of the non-spiking data, which is near zero. To avoid forcing the data to match multiple periods, we fit all 16 models to the first two spikes of the training data (Figure 3A). We then evaluate the fit of the models using the $l^{2}$ norm and Akaike information criteria on the first two spikes of the training data, and validate the models on out-of-sample data, which includes three spikes, using dynamic time warping.
@@ 2.4 Discovered model is consistent with Hourglass and Threshold Hypothesis
The dynamics of the hidden physiological state, $y$, of Model 6 appear with a constant period rather than a period that varies with body temperature (Figure 3C). This supports the Hourglass and Threshold Hypothesis over the Torpor Arousal Clock Hypothesis. The ability to differentiate between the two hypotheses is not dependent on the parameterization and structure of the model; the dimensionality of the thermophysiological system alone supports a constant period and the depleted metabolite hypothesis. Planar ODE systems cannot give rise to non-trivial solutions in which $x(t)$ and $y(t)$ have different periods [38]. However, we can study the identified model’s structure, parameterization, and symmetries to extract additional physiological implications.
Directly relating the hidden state $y(t)$ to a particular metabolite or physiological state is impossible without direct measurements. However, we can computationally explore possible mechanisms and hypothesize about the minimal mechanism of regulation. In the simplest case, we can imagine the hidden state as a molecule that is depleting over time as envisioned by Ruf, Giroud, and Geiser [8] (see Figure 3C top). However, we believe it is important to consider all solutions consistent with the data and their interpretations. By introducing a symmetry transformation into the hidden state of the form $\bar{y}=-y$, which does not change the sparsity pattern of the model, we can generate dynamics in the hidden state that become negative (see Figure 3C bottom), potentially corresponding to the difference between multiple molecules rather than an individual entity.  [^4] Introducing this change of variables to the system results in the the following system $\dot{x}$ $=\theta_{1}x^{2}-\theta_{2}x^{3}+\theta_{3}x\bar{y}-\theta_{4}x^{2}\bar{y}+\theta_{0},$ $x(0)=x_{0},$ $\dot{\bar{y}}$ $=-\theta_{5}x^{2}+\theta_{6}\bar{y}^{2},$ $\bar{y}(0)=\bar{y}_{0}.$ In this case, it is possible for one of the molecules to increase, as may be expected of a toxic molecule that accumulates during torpor. This alternative parameterization of the model also reflects a fact mentioned in [8]: there is no way to distinguish between an accumulating metabolite and a depleting metabolite as the driver of interbout arousals. These symmetry transformations do not affect the fit to the measured state, so we cannot distinguish which of these model variants and corresponding metabolic scenarios are correct. Whether representing a single molecule or interactions between multiple molecules, the effective state generates hourglass-like dynamics.